OpenStudy (anonymous):

Use the substitution method to solve this system x-2y=-1, x+y=5

OpenStudy (agreene):

x-2y=-1 x=-1-2y <=plug in to 2nd equation x+y=5 (-1-2y)+y=5 -y=6 y=6 <= first answer, plug back in to 1st eqn x-2y=-1 x-2(6)=-1 x-12=-1 x=11 <=second answer answers: y=6 x=11